OpenStudy (anonymous):

what is the solution of the system. use substitution. 3x + 2y = 7 -3x + 11 = y a. 6,-3 b. 6,-7 c. -4, 19/2 d. 5,-4

OpenStudy (anonymous):

Solve for one letter... 3x + 2y = 7 3x = 7-2y x=(7-2y)/3 now plug in x into the other equation and solve for x

OpenStudy (anonymous):

it is D. 5,-4?

OpenStudy (anonymous):

Yes.